In an ore containing uranium the ratio of U238 to Pb206 is 3. Calculate the age of ore assuming that all lead present in ore is final stable product of U238 .Take the half-life of U238 to be 4.5x10^9yr     this was answered by your group but how a/a-x became 4/3 when they said a/a-x is 3
 

Grade:12th pass

1 Answers

Arun
5 years ago
Please find the answer to your question
 
Number of atoms of 238U initially / Numbarof atoms of 238U finally = 4 / 3 = a / (a - x)
[∵ Initially one part lead is present with three parts Uranium]
∴ t = 2.303 / λ log α / (α – x )= 2.303 x 4. 5 x 109 / 0.693 log 4/3
= 1.868 x 109 years.
 
Hope it helps now